Find all second order partial derivatives of f. And verify whether fxy=fyx at any (x,y) for each f. F(x,y)=y cos x-e×


1
Expert's answer
2021-03-07T10:02:06-0500

F(x,y)=y*cos(x) - ex

Fx = -y*sin(x) - ex

Fx,x = - y*cos(x) - ex

Fy,x = -sin(x)

Fx,y = -sin(x)

Fy,y =0

Fy = cos(x)